OpenStudy (anonymous):

What is the line of symmetry for the graph of y = -3x^2 + 12x - 11? 1.) x=-4 2.) x=-2 3.) x=2

OpenStudy (luigi0210):

Isn't that where the vertex is at?

OpenStudy (luigi0210):

Solving it calculus wise.. \[\large -3x^2+12x-11\] \[\large y'=-6x+12\] \[\large 0=-6x+12\] \[\large x=2\]

OpenStudy (anonymous):

It's algebra

OpenStudy (luigi0210):

Vertex formula \(\large \frac{-b}{2a}\)

OpenStudy (luigi0210):

So.. \[\LARGE \frac{-12}{2(-3)}=\frac{-12}{-6}=?\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

So x=2 ..?

OpenStudy (luigi0210):

Yup~